### v4

# A Systematic Survey on Large Language Models for Algorithm Design

Abstract.
Algorithm design is crucial for effective problem-solving across various domains. The advent of Large Language Models (LLMs) has notably enhanced the automation and innovation within this field, offering new perspectives and promising solutions. In just a few years, this integration has yielded remarkable progress in areas ranging from combinatorial optimization to scientific discovery. Despite this rapid expansion, a holistic understanding of the field is hindered by the lack of a systematic review, as existing surveys either remain limited to narrow sub-fields or with different objectives. This paper seeks to provide a systematic review of algorithm design with LLMs. We introduce a taxonomy that categorises the roles of LLMs as optimizers, predictors, extractors and designers, analyzing the progress, advantages, and limitations within each category. We further synthesize literature across the three phases of the algorithm design pipeline and across diverse algorithmic applications that define the current landscape. Finally, we outline key open challenges and opportunities to guide future research.
Algorithm design, Large language model, LLM4AD, Optimization, Heuristic, Evolutionary algorithm.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2025/12copyright: noneccs: Computing methodologiesccs: Computing methodologies Search methodologiesccs: Computing methodologies Heuristic function constructionccs: Theory of computation Design and analysis of algorithmsccs: Theory of computation Algorithm design techniques
1. Introduction
Algorithms, which are a sequence of computational steps for solving a well-specified computational problem (cormen2022introduction), play a crucial role in addressing various problems across various domains such as industry, economics, healthcare, and technology (kleinberg2006algorithm; cormen2022introduction). Traditionally, designing algorithm has been a labor-intensive process that demands deep expertise. Recently, there has been a surge in interest towards employing learning and computational intelligence methods techniques to enhance and automate the algorithm development process (bengio2021machine; tang2024learn).
Over the past few years, the application of Large Language Models for Algorithm Design (LLM4AD) has merged as a promising research area with the potential to fundamentally transform the ways in which algorithms are designed, implemented, and optimized. The remarkable capability and flexibility of LLMs have demonstrated potential in enhancing the algorithm design process from algorithm ideation (girotra2023ideas) to implementation (liu2024evolution). This approach not only reduces the human effort required in the design phase but also enhances the creativity and efficiency of the produced solutions (romera2024mathematical; liu2024llm4ad; novikov2025alphaevolve).
Despite this surge of interest, the field lacks a systematic survey with a clear organizational structure. While several recent surveys have reviewed the interaction of LLM and algorithms, they either focus on narrow sub-fields or adjacent but different domains. For example, code generation surveys (wang2023review; zheng2023survey; jiang2024survey; wang2024survey) emphasize translating specifications into executable code, focusing on implementation rather than the upstream ideation and strategic design of algorithms. Reviews on optimization and evolutionary computation (wu2024evolutionary; fan2024artificial; zhang2025systematic; huang2024large) target specific problem classes (e.g., combinatorial optimization), offering limited breadth beyond the target sub-fields. Surveys on LLM agents and planning (pallagani2024prospects; wang2024survey; guo2024large; ferrag2025llm) center on system architecture, tool use, and reasoning pipelines rather than the creation and refinement of core algorithmic logic. As summarized in Table 1, these surveys cover partial stages of algorithm design or specific algorithm types.
This paper aims to fill this gap by providing a systematic survey dedicated to the emerging field of LLM4AD. To ensure conceptual clarity, we first establish a precise definition for “algorithm design” and the scope of this survey, distinguishing it from general-purpose programming. We then introduce a role-based taxonomy that organizes the literature according to the four fundamental ways LLMs are being employed in the algorithm design process. By synthesizing insights from over 180 recent papers, we analyze the using of LLM in different algorithm design stages and application domains. Finally, we critically assess the field’s current challenges and identify promising opportunities for future research. We intend for this survey to serve as an essential resource for both newcomers seeking a structured overview and experts looking for a consolidated analysis of the latest advancements.
Table 1. A Comparison of Different Survey Papers Across Algorithmic Design Stages and Algorithm Types.
Year Survey LLM Algorithmic Design Stages Algorithm Type
Algorithm
Ideation
Algorithm
Implementation
Algorithm
Optimization
Non-heuristic
Heuristic and
Metaheuristic
Hybrid
2023 Wang et al. (wang2023review) Yes No Yes Partial No No No
Zhao et al. (zhao2023automated) Yes Yes Yes Partial No Yes Partial
Zheng et al. (zheng2023survey) Yes No Yes Partial Partial Partial Partial
2024 Ahn et al. (ahn2024large) Yes Partial No No Yes No No
Guo et al. (guo2024large) Yes No Partial Partial Partial Partial Partial
Jiang et al. (jiang2024survey) Yes No Yes Partial No No No
Joel et al. (joel2024survey) Yes No Yes Partial No No No
Wu et al. (wu2024evolutionary) Yes Partial Partial Partial No Partial Partial
Fan et al. (fan2024artificial) Partial No Yes Yes Yes Yes Yes
2025 Ferrag et al. (ferrag2025llm) Yes No Partial Partial Partial Partial Partial
Xiao et al. (xiao2025survey) Yes No No No No No No
Zhang et al. (zhang2025systematic) Yes Partial Partial Partial No Partial Partial
Ma et al. (ma2025toward) Partial Partial Yes Yes No Partial Partial
Da Ros et al. (da2025large) Yes Partial Yes Yes No Partial Partial
Ours Yes Yes Yes Yes Yes Yes Yes
Figure 1. Overview of the Survey Structure: Methodology, Taxonomy of LLM Roles, Stages of LLM‑assisted Algorithm Design, Applications, Open Challenges and Future Directions.
Fig. 1 provides an overview of this survey’s structure. The remainder of the paper is organized as follows. Section 2 outlines our survey methodology, including scope and literature collection and screening pipeline. Section 3 introduces a taxonomy for organizing LLM4AD research, categorizing works based on the primary role of the LLM in the algorithm design process: as optimizer (LLMaO), predictor (LLMaP), extractor (LLMaE), or designer (LLMaD). Subsequently, Section 4 reviews works across different algorithm development stages, while Section 5 summarizes key application domains. Section 6 discusses current open challenges and promising future research directions. Finally, Section 7 presents the conclusions.
2. Methodology
2.1. Definition and Scope
This paper focuses on studies where LLMs substantively contribute to the conception, synthesis, or refinement of algorithms. This section defines our core concepts and delineates the scope of this survey to clarify its boundaries with respect to related research works.
Algorithm
We adopt the definition of an “algorithm” as a well-defined computational procedure that transforms a set of inputs into a set of outputs in a finite amount of time (cormen2022introduction). This includes deterministic and stochastic procedures, exact methods and approximations (kleinberg2006algorithm), as well as heuristics and metaheuristics (glover2006handbook). Examples include a procedure that sorts a set of integers, a method that finds a shortest path on a graph, or a heuristic for a scheduling problem. This scope covers both classic textbook algorithms and practically motivated strategies that trade optimality for speed or simplicity.
Algorithm Design
We cover three stages of algorithm design: i) generation of algorithmic ideas or pseudocodes (si2024can); ii) algorithm implementation, which concerns the production of executable code from specifications (van2025code); and iii) algorithm evaluation, where LLMs are used for assessing the performance and analyzing the behavior of the designed algorithms (liu2024llm4ad). A crucial distinction is made between algorithm design and general-purpose code generation (jiang2024survey). Studies that merely translate algorithms or procedures into code are excluded.
Large Language Models
Our focus is on large-scale language models, typically with billions of parameters, capable of sophisticated text and code processing (zhao2023survey). This includes both text-only and multi-modal LLMs where language is a core component. We exclude smaller-scale models and traditional machine learning approaches for algorithm generation, which have been discussed by other survey papers (bengio2021machine; ma2025toward).
2.2. Paper Collection and Scanning
We introduce the detailed pipeline for paper collection and scanning, which consists of three stages:
  • Stage I Data Extraction and Collection: We collect the related papers through Google Scholar, Web of Science, and Scopus. The logic of our search is the title must include any combinations of at least one of the following two groups of words “LLM”, “LLMs”, “Large Language Model”, “Large Language Models” and “Algorithm”, “Heuristic”, “Search”, “Optimization”, “Optimizer”, “Design”, “Function” (e.g., LLM and optimization, LLMs and algorithm). After removing duplicate papers, we ended up with around 3,000 papers as of October 1, 2025.
    Refer to caption Four stages for paper collection.
    Figure 2. A Three-stage Pipeline for Paper Collection and Screening.
    (a) Annual Publication Volume of Surveyed Papers (until Sep. 2025).
    (b) Distribution on Research Paradigms
    Figure 3. Publication Trends and Paradigm Distribution in the Surveyed Literature.
  • Stage II: Paper Screening This stage involved a two-step screening process to identify the most relevant papers. First, we screened the titles and abstracts of the 3,000 papers against predefined exclusion criteria: i) The paper is not written in English. ii) The paper’s primary focus is not on algorithm design (e.g., it focuses only on general code generation without any algorithmic component). iii) The paper does not utilize large language models as defined in our scope. This initial screening narrowed the corpus down to about 500 papers. Subsequently, we conducted a full-text review of these manuscripts, applying the same exclusion criteria more rigorously to filter out papers that, upon closer inspection, lacked substantive content on LLM-aided algorithm design. This thorough review resulted in a refined set of 150 high-quality, relevant papers.
  • Stage III: Snowballing and Cross-checking To ensure comprehensive coverage and mitigate the limitations of keyword-based searches, we performed a backward snowballing procedure on the 150 papers. This involved manually examining the reference lists of these papers to identify relevant studies that our initial search may have missed (for instance, papers using terminology like “code generation” instead of “algorithm design” but still involving some algorithm design tasks). Additionally, we manually appended a small number of works based on the authors’ domain knowledge to avoid omitting any important contributions. After integrating these additional papers identified through snowballing and expert knowledge, our final corpus contained over 180 papers.
We acknowledge that, given the vast domains of algorithm design and the volume of literature, it is impossible to guarantee an exhaustive coverage of all relevant papers. Instead, our objective is to systematically survey the landscape, focusing on a representative body of work that allows us to organize and discuss the field. Fig. 3(a) illustrates the number of publications per year surveyed in this paper. The graph shows a marked rise in research activity related to LLM4AD, and most of the related studies have been conducted in the last two years.
3. Taxonomy of LLM Roles in Algorithm Design
According to the roles of LLM in algorithm design, existing works can be categorized into four paradigms: LLM as Optimizer (LLMaO), LLM as Predictor (LLMaP), LLM as Extractor (LLMaE), and LLM as Designer (LLMaD). Fig. 3(b) displays the distribution of publications across these four paradigms. This section discusses the progress, advantages, and limitations of each category.
3.1. LLM as Optimizer (LLMaO)
In LLMaO (Fig. 4), LLMs are employed as a black-box optimizer within an algorithmic framework to generate and refine solutions. Fig. 4 (b) illustrates an example applied to the Traveling Salesman Problem (TSP), which involves optimizing a tour that visits each city (node) exactly once and returns to the starting city, with the goal of minimizing the total route length. In this example, we are given four nodes (A, B, C, and D) with their coordinates, and the algorithm has two existing tours (tour 1 and tour 2) each with a different sequence and route length. The LLM is used to generate a potentially better tour with a shorter length. The LLM directly outputs the optimized tour sequence, which is commonly the role of hard-coded optimizers within traditional algorithmic approaches.
Figure 4. (a) Large Language Models as Optimizers (LLMaO). LLMs Serve as Optimizers within the Algorithm to Generate New Solutions. (b) An Example for LLMaO on Traveling Salesman Problem.
One of the initial efforts to utilize LLM as Optimizer in algorithm design is by yang2024large. They leverage the in-context learning capabilities of LLMs to generate new solutions for specific problems based on previously evaluated solutions. This method is applied iteratively to refine solutions further. yang2024large have successfully demonstrated this technique across various domains, including continuous and combinatorial optimization, as well as machine learning tasks.
From an evolutionary algorithm perspective, using LLMs to generate solutions from existing data can be seen as analogous to search operators in EA. For instance, liu2025large introduce the use of LLMs as evolutionary operators to tackle multi-objective problems. This method involves breaking down a multi-objective problem into simpler single-objective tasks, with LLMs acting as black-box search operators for each sub-problem to suggest new solutions. In a related study, liu2024largelmea explore the integration of LLMs within EAs, not just for generating solutions but also for guiding selection, crossover, and mutation processes. Meanwhile, brahmachary2024large propose a new population-based evolutionary framework that includes both exploration and exploitation pools, with solutions being exchanged during the optimization process and LLMs generating solutions for both pools.
Differing from direct solution generation, lange2024large investigate the use of LLMs in designing evolution strategies, introducing a new prompting strategy to enhance the mean statistic in their EvoLLM method, which shows superior performance over baseline algorithms in synthetic black-box optimization functions and neuroevolution tasks. They also demonstrate that fine-tuning LLMs with data from teacher algorithms can further improve the performance of EvoLLM. custode2024investigation present a preliminary study that uses LLMs to automate hyperparameter selection by analyzing optimization logs and providing real-time recommendations. Moreover, xu2025autoep adopt LLMs to adaptively adjust the hyperparameter for metaheuristic algorithms.
Beyond traditional optimization tasks, LLMaO has been widely adopted in prompt engineering for LLMs, a process often referred to as “automatic prompt optimization” (zhou2022large). These methods primarily involve iterative refinement of prompts by LLMs to improve their effectiveness for specific models (typically LLMs). Techniques include resampling-based strategies, where LLMs generate variations of original prompts while maintaining semantic similarity (wang2023promptagent), and reflection-based strategies, where LLMs optimize by analyzing and learning from previous prompt iterations or errors (guo2023connecting), have been explored. ma2024large note that LLM optimizers often struggle to accurately identify the root causes of errors during the reflection process, influenced by their pre-existing knowledge rather than an objective analysis of mistakes. To address these issues, they propose a new approach termed “automatic behavior optimization”, aimed at directly and more effectively controlling the behavior of target models. liu2025language introduce RSBench, a benchmark set specifically for the task of evaluating LLM-based evolutionary algorithms in optimizing recommendation prompts in recommender systems.
Discussion:
Traditional optimizers rely on numerical or symbolic update rules. LLMaO introduces a language-conditioned optimization process, where LLMs propose and refine candidate solutions by reasoning over problem descriptions, instance context, and past trajectories rather than following fixed mathematical updates.
  • Advantages: 1) Leverages pre-trained domain knowledge and natural-language context. For example, in the TSP, conventional operators (e.g., 2-opt (laporte1992traveling)) improve routes based on hard-coded heuristics that do not exploit textual instance information. In contrast, LLM-based optimizers can condition on problem descriptions, constraints, previously evaluated tours and history information to produce informed improvements. liu2024largelmea integrate LLMs as evolutionary optimizers; given task descriptions, existing solutions, and stepwise traces, the LLM performs multi-step search and generates new candidates. 2) Enables adaptive refinement without handcrafted rules. yang2024large show that LLMs can infer optimization directions from prompt-supplied trajectories on small-scale problems, and nie2024importance demonstrate that natural-language feedback during search further enhances LLM-driven optimization.
  • Limitations: 1) Limited interpretability and theoretical guarantees due to the black-box nature. Early efforts approximate LLM behavior with linear operators (liu2025large) and analyze convergence under restricted settings (lee2025convergence), but general frameworks remain open. 2) Sensitivity to prompts and domain priors. Competitive results often require careful prompt design and alignment with the LLM’s training distribution (yang2024large). 3) Computational and token cost at scale. Most evaluations focus on small instances (liu2024largelmea; yang2024large); scaling to large problems is challenging due to long inputs/outputs, inference latency, and costs (liu2025large).
3.2. LLM as Predictor (LLMaP)
LLMaP utilizes LLMs as surrogate models (Fig. 5) to predict the outcomes or responses of solutions, operating in either a classification or regression context (hao2024large). Fig. 5 (b) provides an example on a simple TSP instance, where the LLM is given the task description and a set of existing evaluated tours along with their corresponding lengths. The LLM is then instructed to evaluate a new tour, predicting its length and performance (whether it improves upon the existing tours). Note that while the length of a TSP tour is easily calculated and serves here as a straightforward illustration, LLMaP is designed for tasks where evaluations are typically expensive or difficult to obtain.
Figure 5. (a) Large Language Models as Predictors (LLMaP). LLMs are Utilized Iteratively in Algorithms to Predict a Solution’s Outcomes or Responses. (b) An Example for LLMaP on Traveling Salesman Problem.
The majority of LLMaP works use LLMs as pre-trained models as a regression model to predict solution scores. For instance, LLMs have been used as performance predictors for deep neural network architectures by jawahar2024llm. It offers a cost-effective alternative for performance estimation in neural architecture search. zhang2024can introduce LINVIT, an algorithm that incorporates guidance from LLMs as a regularization factor in value-based RL to improve sample efficiency. Science discovery is another domain that LLMaP has commonly investigated. For example, li2023codonbert introduce CodonBERT for sequence optimization of mRNA-based vaccines and therapeutics. CodonBERT uses codons as inputs and is trained on over 10 million mRNA sequences from various organisms. soares2024capturing demonstrate the use of LLMs in predicting the performance of battery electrolytes. Other applications include employing LLMs to determine the fame score of celebrities to predict the box office performance of projects in the motion pictures industry (alipour2024data) and adopting LLMs to score the video question answering by using detailed video captions as content proxies (zhang2024direct).
For classification, hao2024large introduce LAEA, which employs LLMs as surrogate models within evolutionary algorithms for both regression and classification, eliminating the need for costly model training. In another study, chen2023label develope a label-free node classification method that leverages LLMs to annotate nodes. These annotations are subsequently used to train graph neural networks, resulting in enhanced performance. Moving beyond binary classification, bhambri2024efficient utilize LLMs to predict discrete actions for constructing reward shaping functions in Reinforcement Learning (RL). Their method demonstrate effectiveness within the BabyAI environment, showcasing the versatility of LLMs in various settings. wang2024resllm explore the use of LLMs in federated search, applying them in a zero-shot setting to effectively select resources. This approach highlights the potential of LLMs in improving resource selection without prior explicit training on specific tasks.
Discussion:
Conventional surrogate models (e.g., Gaussian processes (GPs) or neural regressors (jin2018data)) rely on structured features and explicit training. In contrast, LLMaP leverages pre-trained LLMs as semantic surrogates that can interpret textual and multimodal context when predicting outcomes.
  • Advantages: 1) Effective in data-scarce or concept-driven tasks due to embedded general knowledge (egami2024using). For example, wong2024generative employ a multimodal LLM to score car shapes, accelerating early-stage design by filtering out poor candidates before costly simulation signals that conventional surrogates built on numeric features struggle to capture. ge2025mora adopt LLMs to extract deeper interest preferences from the user’s behaviour and interaction history to dynamically adjust the prediction of user’s rating of items in the recommendation algorithm. 2) Requires little or no retraining, reducing computational cost. As discussed by hao2024large, conventional surrogates often require repeated rebuilding or fine-tuning as new solutions are sampled, which is expensive for large-scale cases. LLMs can be used zero-shot or with lightweight in-context examples and can use flexible descriptors (text, code, images), thereby avoiding tedious feature engineering.
  • Limitations: 1) Quantitative precision is often inferior to specialized regression models. In the evaluation by hao2024large, LLMs, though flexible, often achieve lower accuracy compared to GP models trained directly on the target data. xie2025large propose hybrid schemes that combine LLMs with conventional surrogates: instead of predicting scores directly, the LLM selects which specialized surrogate to use, offering a pragmatic compromise between performance and efficiency. 2) Outputs are sensitive to prompt design and phrasing. For example, taboada2025ontology use LLMs for ontology alignment and find that simple prompts limit performance, whereas incorporating contextual ontology information improves matching accuracy. This sensitively introduces additional costs for prompt refinement, which relies on domain knowledge.
3.3. LLM as Extractor (LLMaE)
LLMaE leverages LLMs to mine and extract embedded features or specific knowledge from target problems and/or algorithms, which are then used to enhance algorithm-based problem solving (Fig. 6). As shown in Fig. 6 (b), when provided with the task description and tours for a TSP instance, LLMs can be instructed to extract key features, such as node count, which can then be utilized in algorithm design. These features are typically defined by experts or learned implicitly using conventional machine learning models.
Figure 6. (a) Large Language Models as Extractors (LLMaE). LLMs are Employed to Extract Features or Specific Knowledge from Target Problem and/or Algorithms. (b) An Example for LLMaE on Traveling Salesman Problem.
Beyond embedding-based feature extraction, LLMs excel in text comprehension and knowledge extraction, allowing them to discern subtle patterns and relationships within the data that might not be evident through conventional feature extraction methods. For example, wu2024large utilize LLMs to extract high-dimensional algorithm representations by comprehending code text. These representations are combined with problem representations to determine the most suitable algorithm for a specific problem. du2024mixture propose a mixture-of-experts framework augmented with LLMs to optimize various wireless user tasks. The LLM is used to analyze user objectives and constraints, thus selecting specialized experts, and weighing decisions from the experts, reducing the need for training new models for each unique optimization problem. Additionally, memduhouglu2024enriching use LLM to enhance the classification of urban building functions by interpreting OpenStreetMap tags and integrating them with physical and spatial metrics. Traditional techniques, which have previously struggled with semantic ambiguities, are outperformed by LLMs due to their superior ability to capture broader language contexts.
Typical feature extraction relies on statistical learning or dimensionality reduction from structured data. LLMaE instead is able to perform semantic extraction, deriving contextual features and domain knowledge from unstructured sources such as text, code, or documentation.
  • Advantages: 1) It generates concept-aware embeddings that integrate both linguistic and symbolic meaning. For instance, xu2025evospeak use LLMs to analyze existing heuristic structures and extract underlying design principles and domain-relevant insights, thereby providing a warm start to enhance the quality of designed heuristics. 2) It combines textual, spatial, and numeric features into a unified representation. For example, (wu2024large) use LLMs to extract algorithm features for algorithm selection. Unlike traditional approaches, such as machine learning prediction models, both text descriptions and code implementations can be leveraged.
  • Limitations: 1) There is a risk of hallucinated or misinterpreted features in specialized domains. For instance, (hu2025discovering) report information loss and misinterpretation when using multi-modal LLMs for designing algorithms for agents. 2) The validation and interpretability of embeddings remain challenging. jiang2024bridging use LLM embeddings to solve the vehicle routing problem. While the extracted information proves beneficial in ablation studies, providing a clear interpretation and analysis of the embedding remains difficult.
3.4. LLM as Designer (LLMaD)
LLMaD directly creates algorithms or specific components (Fig. 7). LLMs can generate heuristics (liu2024evolution), write code snippets (hemberg2024evolving), and formulate functions (romera2024mathematical) that integrate seamlessly into algorithmic systems, or even design the entire algorithms (van2024llamea; novikov2025alphaevolve). Fig. 7 demonstrates this process using a TSP instance. Unlike other paradigms, LLMaD directly produces new algorithms or algorithm components based on a task description and an existing algorithm implementation.
Figure 7. (a) Large Language Models as Designers (LLMaD). LLMs are Used to Directly Create the Entire Algorithms or Specific Components. (b) An Example for LLMaD on Traveling Salesman Problem.
Function design is among the early applications of LLMaD. Eureka (ma2024eureka) leverages the capabilities of LLMs in code-writing, and in-context learning to evolve and optimize reward functions for RL. It can generate reward functions without specific prompts or predefined templates, achieving better performance than rewards designed by human experts. Similarly, Auto MC-Reward (li2024auto) utilizes LLMs to automatically design dense reward functions for RL agents in environments with sparse rewards. The three key components of Auto MC-Reward work together to iteratively refine the reward function based on feedback from the agent’s interactions with the environment. Through this iterative process, the agent is able to learn complex tasks more efficiently, as demonstrated in experiments in Minecraft. Moreover, FunSearch (romera2024mathematical) adopts LLMs for function generation in an evolutionary framework with a multi-island population management. It demonstrates promising results on both mathematical problems and combinatorial optimization problems.
EoH (liu2024evolution; liu2023algorithm) presents an early attempt to adopt the LLM as a designer for Automated Heuristic Design (AHD). It uses both heuristic ideas and code implementations to represent heuristics and adopts LLM in an evolutionary framework to create, combine, and revise the heuristics. While it is original proposed for heuristic design, it has been applied on different algorithm design tasks including combinatorial optimization problems (liu2024evolution; yao2024multi; liu2025eoh), Bayesian optimization (yao2024evolve), image adversary attack (guo2024autoda), and edge server task scheduling (wang2024ts), among others. Moreover, LLaMEA (van2024llamea; van2024loop) develops an iterative framework to generate, mutate, and select algorithms based on performance metrics and runtime evaluations. The automatically designed algorithms outperform state-of-the-art optimization algorithms on some benchmark instances. ReEvo (ye2024reevo) introduces an evolutionary framework with both short and long-term reflections, which provides a search direction to explore the heuristic space. In addition to evolutionary search framework, recent attempts also adopt other frameworks such as large neighborhood search (xie2025llm) and Monte Carlo Tree Search (MCTS) (zhengmonte) for effective explore the algorithm space. Unlike previous studies that focus on optimizing a single performance criterion, MEoH (yao2024multi) considers multiple performance metrics, including optimality and efficiency, and seeks a set of trade-off algorithms in a single run in a multi-objective evolutionary framework. A dominance-dissimilarity score is designed for effectively searching the complex algorithm space.
LLM-based agent design has also gained much attention. For example, ADAS (hu2024automated) proposes an automated design of agentic systems, which aims to automatically generate powerful agentic system designs by using meta agents that program new agents. They present a novel algorithm, meta agent search, which iteratively creates new agents from an archive of previous designs, demonstrating through experiments that these agents can outperform state-of-the-art hand-designed agents across various domains. Further studies on LLM-based agent systems are discussed in (sun2024llm) and (liu2024agentlite).
Discussion:
Different from existing approaches, such as human-crafted heuristics, AutoML, and genetic programming, search within explicit procedural representations. LLMaD advances this by using LLMs for language-based algorithm synthesis, generating algorithms or components via natural-language reasoning and code generation.
  • Advantages: 1) Accelerates algorithm design by reducing human intervention and implementation effort. LLMaD often operates zero-shot or with lightweight iterative prompting; the outputs are executable text/code that can be inspected, profiled, and reused, while achieving competitive efficiency (liu2024evolution). 2) Supports iterative self-improvement when embedded in feedback or evolutionary loops, expanding the creativity and diversity of algorithmic solutions (zhang2025darwin; van2025code). For instance, zhang2025darwin propose a tree-based search framework that uses LLMs to design diverse, high-quality agents in an open-ended manner.
  • Limitations: 1) LLMs struggle with synthesizing complete, state-of-the-art algorithms for complex tasks; most studies target specific components (e.g., heuristics, reward functions, code snippets) rather than the complete solvers (novikov2025alphaevolve). Challenges around correctness, robustness, and runtime guarantees persist. 2) Algorithm design is domain-specific, and standalone LLMs typically lack sufficient task-specific knowledge, leading to subpar performance (zhang2024understanding). Effective LLMaD systems therefore depend on search frameworks that iteratively test, execute, and refine designs while interacting with the environment and tools (zhang2024understanding; van2025blade).
Figure 8. The using of LLMs in Different Stages of Algorithm Design: 1) Algorithm Ideation, 2) Algorithm Implementation, and 3) Algorithm Evaluation.
4. Stages of LLM‑Assisted Algorithm Design
Algorithm design is a multi-stage process where LLMs can provide targeted assistance at distinct phases. We identify three key stages in this process: 1) algorithm ideation, 2) algorithm implementation, and 3) algorithm evaluation. This section organizes the existing literature along these three stages, summarizing the typical artifacts, recent progresses and challenges.
4.1. Stage I: Algorithm Ideation
The ideation stage encompasses both problem formulation and the generation of novel algorithmic concepts.
Problem Formulation.
While not a direct step in creating an algorithm, modeling the target problem is a crucial precursor to its design, and LLMs are increasingly used to automate this process. Early systems, such as OptiMUS, guide users from natural language descriptions to formal linear programs, producing executable solver (ahmaditeshnizi2024optimus). To mitigate the privacy and controllability concerns associated with proprietary APIs, ORLM combines an instruction-following recipe (OR-Instruct) with open-source backbone models, reporting strong results on benchmarks like NL4Opt and the new IndustryOR benchmark (tang2024orlm; ramamonjison2023nl4opt). Complementary fine-tuning strategies, such as those in LM4OPT, demonstrate that even modest-sized models can be specialized to reliably map domain-specific text to optimization modeling (ahmed2024lm4opt). In the context of interactive decision support, DOCP integrates user feedback to progressively assemble problem-specific models before invoking solvers (wasserkrug2024large). For a comprehensive overview, xiao2025survey provide an in-depth survey of LLMs for optimization modeling.
Algorithm Concept Design.
LLMs have shown promise in generating new algorithmic ideas across various domains (si2024can). For instance, pluhacek2023leveraging use LLMs to analyze and decompose well-performing algorithms to propose hybrid algorithms that combine their complementary strengths. While this work demonstrated that novel algorithmic ideas could be designed, it did not implement them for specific algorithm design tasks. EoH presents an early attempt to co-design algorithm concepts in natural language alongside their executable code implementations (liu2024evolution). Different strategies have been explored to enhance this co-design search process, including reasoning over algorithmic ideas during design (bomer2025leveraging; ling2025complex) and controlling the diversity to foster innovation (yao2025evolution).
However, how to effectively measure the quality and novelty of an algorithmic idea remains an open question (liu2025fitness). Moreover, the concept descriptions are often highly abstract, which lacks a strict mapping to a detailed implementation, therefore making it hard to effectively guide algorithm implementation (gurkan2025lear).
4.2. Stage II: Algorithm Implementation
The implementation stage spans from initial code synthesis to subsequent performance optimization.
Code Generation.
Given an algorithmic idea or a natural language description, LLMs can be instructed to generate the corresponding code implementation (liu2024evolution; yao2024evolve; zhengmonte). Some works bypass the explicit ideation stage and directly use code as the primary representation of an algorithm (romera2024mathematical; ye2024reevo). Instead of generating code from structure, ADAS employs meta-agents to program new agents from an archive of prior designs, thereby automating the construction of multi-component pipelines (hu2024automated). Initially, code generation focused on individual algorithmic components, reflecting the limited capabilities of early LLMs and simpler design pipelines (romera2024mathematical; liu2024evolution). More recently, this has been extended to generating entire algorithms (van2024llamea). To handle large-scale code, novikov2025alphaevolve use “diff blocks” to highlight specific sections, enabling the model to generate or revise partial code instead of regenerating the entire program. To improve reliability at scale, LLaMoCo introduces instruction tuning tailored to code generation for optimization tasks, reducing the reliance on expert-crafted prompts (ma2024llamoco).
Despite these advances, the generation of entire, complex algorithms remains a significant challenge. First, the reliable generation of long programs is a difficult problem in itself (jiang2024survey). Second, it is challenging to ensure that a complicated implementation correctly instantiates the intended algorithmic idea and is effective for the target task. Addressing this requires not only powerful LLMs but also sophisticated frameworks, such as multi-agent systems (Ishibashi2024selforganized), to manage the complexity.
Algorithm Optimization.
A common pattern for algorithmic code optimization is closed-loop refinement, where execution feedback informs subsequent code edits and integration choices (zhang2024understanding). Evolutionary search is the most prevalent framework for this purpose, maintaining a population of algorithms that are progressively refined by an LLM (zhang2024understanding). Variants of this simple evolutionary approach have emerged. For instance, ReEvo combines short- and long-term reflection with information on history evaluated results to guide algorithm optimization (ye2024reevo), while HSEvo adopts diversity control strategies with two population diversity measurements to enhance the optimization process (yao2025evolution). Recently, non-evolutionary search frameworks, such as large neighborhood search and Monte Carlo tree search (zhengmonte), have also been demonstrated to be effective. Moreover, instead of prompting LLMs to optimize algorithm implementations, recent attempts have explored fine-tuning LLMs, in both offline (liu2025fine) and online manners (huang2025calm), to learn a preference for generating better-performing algorithms during algorithm optimization for specific design tasks.
A significant open challenge is ensuring that the optimized algorithms generalize robustly. An algorithm may be overfitted to the specific benchmarks or problem instances used during the optimization loop, failing to perform well on unseen instances (sim2025beyond) or under different conditions (shi2025generalizable).
4.3. Stage III: Algorithm Evaluation
The final stage, algorithm evaluation, involves assessing the performance and analyzing the behavior of the designed algorithms.
Performance Testing.
To ensure comparability and robustness in LLM-driven algorithm design, recent efforts have focused on establishing standardized benchmarks. Notable examples include CO-Bench and HeuriGym, which are tailored for structured combinatorial optimization tasks (sun2025co; chen2025heurigym). In addition to combinatorial optimization, LLM4AD (liu2024llm4ad) provides a unified evaluation platform that spans a broader range of domains, including optimization, machine learning, and scientific discovery (liu2024llm4ad). Beyond static benchmarks for testing, some approaches leverage LLMs to dynamically refine the evaluation process itself. For instance, li2025cocoevo propose a method to co-evolve algorithms and their test instances, aiming to generate more effective test cases that better guide the algorithm design process. Similarly, duan2025ealg introduce a mutation-based adversarial approach that dynamically evolves instance generation procedures to create increasingly difficult problems, thereby enhancing the generalization performance of the designed algorithms.
Moving beyond a single performance score, a deeper evaluation should analyze the algorithm’s behavior, including its search trajectory, robustness across different problem distributions, convergence properties, and computational efficiency (yao2024multi). LLMs are uniquely positioned to automate this in-depth analysis. Just as a human expert would, they can interpret search patterns, identify failure modes, and diagnose the root causes of poor performance to provide actionable, human-like insights to directly enhance the design process (yu2024deep).
Algorithm Analysis.
In addition to generation, LLMs are also being employed to analyze algorithmic outcomes and explain their behavior. d2024exploring empirically analyzed and compared the quality of explanations provided by three different LLMs for seven state-of-the-art quantum algorithms. Their findings indicated that while the explanations were consistent across multiple iterations, there remains a significant gap for improvement in quality, and the results were highly sensitive to the specific prompts used. In a different application, chacon2024large integrated LLMs into STNWeb, an algorithm analysis tool. The LLM generates extensive written reports, complemented by automatically generated plots, which enhances the user experience and lowers the barrier to adoption for the broader research community.
The central challenge for LLM-based algorithm analysis is the gap between descriptive summarization and causal, principled understanding. Current methods can generate consistent and well-written reports, but these often remain surface-level descriptions of behavior rather than deep explanations rooted in algorithmic theory (chacon2024large). The analysis quality is also highly sensitive to prompt engineering, making it unreliable and not yet a trustworthy source of foundational insight. For the field to advance, LLMs must move beyond post-hoc commentary to providing causal reasoning about why an algorithm performs as it does. For instance, by explain its convergence properties, time complexity, or failure modes based on its structural components (kleinberg2006algorithm; van2025code; liu2025fitness).
5. Applications
5.1. Optimization Algorithms
In this subsection, we delve into the applications of LLMs in designing optimization algorithms. We categorize the existing literature into combinatorial optimization and continuous optimization. Then we proceed to compare the various roles played by LLMs, and the specific problems or tasks to which they are applied. The comparative analysis is summarized in Table 2, where we list the names of the frameworks or methods proposed by the authors. For studies that do not explicitly name their methods, we assign appropriate designations in our article and denote them with asterisks for easy reference (e.g., MH-LLM* for (sartori2024metaheuristics)).
5.1.1. Combinatorial Optimization
In the domain of Combinatorial Optimization (CO), automated algorithm heuristics design has been a significant area of interest for a long time. The Traveling Salesman Problem (TSP) stands out as one of the most renowned CO problems, involving the quest for the shortest route to visit all specified locations exactly once and return to the starting point. Some recent work leverages LLMs to evolve algorithms within Evolutionary Computation (EC) framework, such as AEL (liu2023algorithm), ReEvo (ye2024reevo) and EoH (liu2024evolution). Differently, OPRO (yang2024large) employs LLMs as optimizers with a proposed meta-prompt, in which the solution-score pairs with task descriptions are added in each optimization step. Additionally, LMEA (liu2024largelmea) investigates the utilization of LLMs as evolutionary combinatorial optimizers for generating offspring solutions, wherein a self-adaptation mechanism is introduced to balance exploration and exploitation. The Capacitated Vehicle Routing Problem (CVRP) extends the TSP by introducing constraints related to vehicle capacity. To address this challenge, MLLM (huang2024multimodal) devises a multi-modal LLM-based framework with textual and visual inputs to enhance optimization performance. In addition to routing problems, other combinatorial optimization problems that have also been investigated include cap set (romera2024mathematical), bin packing (liu2024evolution), flow shop scheduling (liu2024evolution), hybrid job shop scheduling (li2025llm) and social networks problems (sartori2024metaheuristics).
5.1.2. Continuous Optimization
For single-objective optimization problems with continuous variables, LLaMEA (van2024llamea) utilizes LLMs to automate the evolution of algorithm design. It demonstrates the effectiveness in generating new metaphor-based optimization algorithms on BBOB benchmark (hansen2010black) within IOHexperimenter benchmarking tool (de2024iohexperimenter), which supports evaluating the quality of the generated algorithms and also provides feedback to the LLM during evolution. Instead of creating new algorithms, EvolLLM (lange2024large) introduces a prompt strategy that enables LLM-based optimization to act as an Evolution Strategy (ES) and showcases robust performance on synthetic BBOB functions and neuroevolution tasks. OPRO (yang2024large) illustrates that LLMs can effectively capture optimization directions for linear regression problems by leveraging the past optimization trajectory from the meta-prompt. Additionally, LEO (brahmachary2024large) devises an explore-exploit policy using LLMs for solution generation, the method has been tested in both benchmark functions as well as industrial engineering problems. Different with directly employing LLMs for generating solutions, LAEA (hao2024large) introduces LLM-based surrogate models for both regression and classification tasks and has been validated on 2D test functions using nine mainstream LLMs.
Beyond a single objective, there are multiple competing objectives that need to be optimized simultaneously in many scenarios, forming the multi-objective optimization problems (MOPs). The goal is to identify a set of optimal solutions, referred to as Pareto optimal solution set. An initial exploration of utilizing LLMs to tackle MOPs is introduced in MOEA/D-LMO (liu2025large). Benefiting from the decomposition-based framework, the in-context learning process of LLMs is easily incorporated to generate candidate solutions for each subproblem derived from the original MOP. In the realm of large-scale MOPs, LLM-MOEA* (singh2024enhancing) showcases the inferential capabilities of LLMs in multi-objective sustainable infrastructure planning problem. The study highlights the LLM’s proficiency in filtering crucial decision variables, automatically analyzing the Pareto front, and providing customized inferences based on varying levels of expertise. Additionally, CMOEA-LLM (wang2024large) leverages LLMs with evolutionary search operators to address the challenges of constrained MOPs and exhibits robust competitiveness in DAS-CMOP test suite (fan2020difficulty).
In various real-world applications, the cost of evaluating the objective functions can be very expensive, which greatly limits the evaluation budget in the optimization process (frazier2016bayesian). Bayesian optimization (BO) stands out as a sample-efficient method, it typically employs a surrogate model to approximate the expensive function and well-designed Acquisition Functions (AFs) to carefully select potential solutions. To facilitate the direct generation of solutions using LLMs, HPO-LLM* (zhang2023using) provides LLMs with an initial set of instructions that outlines the specific dataset, model, and hyperparameters to propose recommended hyperparameters for evaluation in Hyperparameter Optimization (HPO) tasks. Furthermore, LLAMBO (liu2024largellambo) incorporates LLM capabilities to enhance BO efficiency, in which three specific enhancements throughout the BO pipeline have been systematically investigated on tasks selected from HPOBench (eggensperger2021hpobench). Instead of utilizing LLMs for direct solution generation, BO-LIFT* (ramos2023bayesian) utilizes predictions with uncertainties provided by a Language-Interfaced Fine-Tuning (LIFT) framework (dinh2022lift) with LLMs to perform BO for catalyst optimization using natural language. EvolCAF (yao2024evolve) introduces a novel paradigm integrating LLMs within EC framework to design AFs automatically for cost-aware BO, the approach showcases remarkable efficiency and discovers novel ideas not previously explored in existing literature on AF design. Similarly, FunBO (aglietti2025funbo) found novel and well-performing AFs for BO by extending FunSearch (romera2024mathematical), the discovered AFs are evaluated on various synthetic and HPO benchmarks in and out of the training distribution.
Table 2. An Overview of Optimization Applications Utilizing Language Models Across Various Domains and Task.
Application Method Role of LLM Specific Problems or Tasks
Combinatorial Optimization EoH (liu2024evolution; liu2023algorithm) LLMaD TSP, Online BPP, FSSP
ReEvo (ye2024reevo) LLMaD TSP, CVRP, OP, MKP, BPP, DPP
OPRO (yang2024large) LLMaO TSP
LMEA (liu2024largelmea) Mixed TSP
MLLM (huang2024multimodal) Mixed CVRP
FunSearch (romera2024mathematical) LLMaD Cap Set Problem, Online BPP
MH-LLM* (sartori2024metaheuristics) LLMaD Social Networks Problem
SolSearch (sheng2025solsearch) LLMaD Satisfiability Problem
LMPSO (shinohara2025large) LLMaO TSP
LLM-NSGA (wan2025surgery) LLMaD Surgery Scheduling Problem
STRCMP (li2025strcmp) LLMaD MILP, SAT
EvoCut (yazdani2025evocut) LLMaD MILP
Continuous Optimization LLaMEA (van2024llamea) LLMaD BBOB
EvoLLM (lange2024large) LLMaO BBOB, Neuroevolution
OPRO (yang2024large) LLMaO Linear Regression
LEO (brahmachary2024large) LLMaO Numerical Benchmarks, Industrial Engineering Problems
LAEA (hao2024large) LLMaP Ellipsoid, Rosenbrock, Ackley, Griewank
MOEA/D-LMO (liu2025large) LLMaO Mulit-objective Synthetic Functions
LLM-MOEA* (singh2024enhancing) LLMaE Multi-objective Sustainable Infrastructure Planning Problem
CMOEA-LLM (wang2024large) LLMaO DAS-CMOP
HPO-LLM* (zhang2023using) LLMaO HPOBench
LLAMBO (liu2024largellambo) Mixed Bayesmark, HPOBench
BO-LIFT* (ramos2023bayesian) LLMaD Catalyst Optimization
EvolCAF (yao2024evolve) LLMaD Synthetic Functions, HPO
FunBO (aglietti2025funbo) LLMaD Synthetic Functions, HPO
LLM-SAEA (xie2025large) LLMaP Synthetic Functions
AwesomeDE (yang2025large) LLMaO Synthetic Functions
BBNM (lee2025convergence) LLMaO Wireless Networks
LLM4CMO (chen2025llm4cmo) LLMaO Constrained Multiobjective Optimization
5.2. Machine Learning Systems and Data Science
In this subsection, we investigate the applications of LLMs in the machine learning domain, focusing on their contribution to algorithmic design. These applications are summarized in Table 3.
5.2.1. Reinforcement Learning
Reinforcement Learning (RL) has been the de facto standard for sequential decision-making tasks, and recently, the synergy between RL and LLMs has emerged as a novel trend in the domain. This convergence mirrors the dynamics of task planning, yet places RL at the core of its methodology. Many of the LLM4AD papers on RL is for automatically designing the reward functions (bhambri2024efficient; ma2024eureka; narin2024evolutionary). In addition, shah2023navigation investigate the employment of LLMs for heuristic planning to steer the search process within RL frameworks. zhang2024can integrate LLMs into RL by introducing a Kullback-Leibler divergence regularization term that aligns LLM-driven policies with RL-derived policies. LLMs have also extended their reach to multi-agent RL scenarios, as shown by du2024mixture, who illustrates their application within a Mixture-of-Experts system to direct RL models in the realm of intelligent network solutions.
5.2.2. Neural Architecture Search
Neural Architecture Search (NAS), which is a significant focus within the AutoML community, has been investigated in many LLM4AD papers. For example, chen2024evoprompting have integrated LLMs with evolutionary search to successfully generate NAS code for diverse tasks. nasir2023llmatic introduce a quality-diversity algorithm tailored to NAS, producing architectures for CIFAR-10 and NAS-bench-201 benchmarks. Moreover, morris2024llm introduce guided evolution for the development of neural architectures and suggest the concept of the evolution of thought in algorithm design. Except for using LLM for design, jawahar2024llm employ LLMs in predicting NAS performance, combining this approach with evolutionary search to effectively create novel network architectures. In contrast to the LLM-based architecture design and performance prediction, zhou2024design explore the adoption of LLMs for transferring design principles to narrow and guide the search space.
5.2.3. Prompt Tuning
Prompt tuning aims to identify the most effective task prompt to enhance the performance of the LLM on a specific task dataset. Despite of requiring specialized training for each specific task, traditional discrete or continuous approaches (li2021prefix) typically necessitate access to the logits or internal states of LLMs, which may not be applicable when the LLM can only be accessed through an API. To address these issues, recent works propose to model the optimization problem in natural language with LLMs as prompts. APE (zhou2022large) utilizes the LLM as an inference model to generate instruction candidates directly based on a small set of demonstrations in the form of input-output pairs. This approach has demonstrated human-level performance on various tasks, including Instruction Induction (honovich2022instruction) and Big-Bench Hard (BBH) (suzgun2022challenging). OPRO (yang2024large) enables the LLM as an optimizer to gradually generate new prompts based on the full optimization trajectory, the optimizer prompt showcases significant improvement compared with human-designed prompts on BBH and GSM8K (cobbe2021training). Inspired by the numerical gradient descent method, APO (pryzant2023automatic) conducts textual “gradient descent” by identifying the current prompts’ flaws and adjusting the prompt in the opposite semantic direction of the gradient. Similar practices are also found in the gradient-inspired LLM-based prompt optimizer named GPO (tang2024unleashing), as well as the collaborative optimization approach (guo2024two) integrating a gradient-based optimizer and an LLM-based optimizer. Differently, guo2023connecting introduce a discrete prompt tuning framework named EvoPrompt that prompts LLM to act like evolutionary operators in generating new candidate prompts, harnessing the benefits of evolutionary algorithms that strike a good balance between exploration and exploitation. StrategyLLM (gao2023strategyllm) integrates four LLM-based agents—strategy generator, executor, optimizer, and evaluator—to collaboratively induce and deduce reasoning. This method generates more generalizable and consistent few-shot prompts than CoT prompting techniques.
5.2.4. Graph Learning
Graph learning is another application with the advancing capabilities of LLMs in symbolic reasoning and graph processing. For example, chen2023label apply LLMs to the task of labeling in Text-Attributed Graphs (TAGs), capitalizing on the language task proficiency of LLMs. Both mao2024identify and chen2024hin adopt LLMs in an evolutionary framework for designing functions. The former evolves heuristic code functions to identify critical nodes in a graph while the latter identifies meta-structures within heterogeneous information networks to enhance their interpretability. Moreover, knowledge graphs have also seen substantial benefits from the application of LLMs. zhou2024autoalign introduce AutoAlign, a method that employs LLMs to semantically align entities across different knowledge graphs, and feng2023knowledge develop the knowledge search language to effectively conduct searches within knowledge graphs.
5.2.5. Dataset Labeling
LLMs have been used for mining semantic and multi-modal information from datasets. LLMs are employed to train interpretable classifiers to extract attributes from images (chiquier2024evolving) and to generate label functions for weakly supervised learning (guan2023can).
Table 3. An Overview of Machine Learning Applications Utilizing Language Models Across Various Domains and Tasks.
Application Method Role of LLM Specific Problems or Tasks
Reinforcement Learning LFG (shah2023navigation) LLMaP ObjectNav Tasks, Real-world Tasks
SLINVIT (zhang2024can) LLMaP ALFWorld Tasks, InterCode Tasks, BlocksWorld Tasks
MEDIC (bhambri2024efficient) LLMaP BabyAI Tasks
Eureka (ma2024eureka) LLMaD IsaacGym Tasks, Bidexterous Manipulation Tasks
EROM (narin2024evolutionary) LLMaD IsaacGym Tasks
LLM-MOE (du2024mixture) LLMaP Intelligent Networks
Neural Architecture Search EvoPrompting (chen2024evoprompting) LLMaD MNIST Dataset, CLRS Algorithmic Reasoning
HS-NAS (jawahar2024llm) LLMaP Machine Translation Tasks
LLMatic (nasir2023llmatic) LLMaD CIFAR-10 Dataset, NAS-bench-201 Benchmarks
LAPT (zhou2024design) LLMaD NAS201, Trans101, DARTs
LLM-GE (morris2024llm) LLMaD CIFAR-10 Dataset
Prompt Tuning APE (zhou2022large) LLMaO Instruction Induction, BBH
OPRO (yang2024large) LLMaO GSM8K, BBH
APO (pryzant2023automatic) LLMaO NLP Benchmark Classification Tasks
GPO (tang2024unleashing) LLMaO Reasoning, Knowledge-intensive, NLP Tasks
MaaO (guo2024two) LLMaO NLU Tasks, Image Classification Tasks
EvoPrompt (guo2023connecting) LLMaO Language Understanding and Generation Tasks, BBH
StrategyLLM (gao2023strategyllm) Mixed Reasoning Tasks
Graph Learning LLM-Critical (mao2024identify) LLMaD Critical Node Identification
LLM-GNN (chen2023label) LLMaP Label-free Node Classification
ReStruct (chen2024hin) LLMaE Meta-structure Discovery
AutoAlign (zhou2024autoalign) LLMaE Entity Type Inference
KSL (feng2023knowledge) LLMaP Knowledge Search
AutoSGNN (mo2025autosgnn) LLMaD GNNs
Dataset Labeling Inter-Classier (chiquier2024evolving) LLMaO iNaturalist Datasets, KikiBouba datasets
DataSculpt (guan2023can) LLMaP Label Function Design
Other Applications LLM2FEA (wong2024llm2fea) LLMaP Objective-oriented Generation
tnGPS (zeng2024tnGPS) LLMaD Tensor Network Structure Search
L-AutoDA (guo2024autoda) LLMaD Adversarial Attack
L-SFE (wang2025llm) LLMaD Causal Structure Learning
5.2.6. Other Applications
Other applications of LLMs extend to a myriad of machine learning tasks. Notably, wong2024llm2fea capitalize on multi-task learning and the iterative refinement of prompts to foster innovative design approaches. zeng2024tnGPS integrate LLMs with evolutionary search to develop a heuristic function aimed at efficiently sifting through candidate tensor sets representing the primal tensor network. Furthermore, guo2024autoda have blazed a trail in employing LLMs to generate novel decision-based adversarial attack algorithms, thus opening up a new diagram for the automatic assessment of model robustness.
5.3. Scientific and Symbolic Domains
This subsection is dedicated to exploring LLM-based scientific discoveries related to algorithm designs. Table. 4 lists the related works in this domain.
5.3.1. Symbolic Regression
In the realm of scientific discovery, LLMs are usually adopted for equation or functioin search. Notably, du2024llm4ed introduce LLM4ED, a framework that employs iterative strategies, including a black-box optimizer and evolutionary operators, to generate and optimize equations. This approach has shown significant advancements in stability and usability for uncovering physical laws from nonlinear dynamic systems. Similarly, shojaee2024llm present LLM-SR, which combines LLMs’ extensive scientific knowledge and code generation capabilities with evolutionary search. This framework excels in proposing and refining initial equation structures based on physical understanding, outperforming traditional symbolic regression methods in discovering physically accurate equations across multiple scientific domains. A bilevel optimization framework, named SGA, is introduce by ma2024llm. It merges the abstract reasoning capabilities of LLMs with the computational power of simulations. This integration facilitates hypothesis generation and discrete reasoning with simulations for experimental feedback and optimization of continuous variables, leading to improved performance.
5.3.2. Chemistry, Biology & Physics
In the field of chemistry, LLMs can be applied not only to conventional molecular generation and design (ishida2024large; bhattacharya2024large), but also to specialized areas such as drug molecule design (ye2024novo), chemical reaction prediction and optimization (rankovic2023bochemian), and catalyst design (wang2024catalm), providing customized solutions. Furthermore, LLMs have also shown promising application prospects in materials discovery (jia2024llmatdesign; zhang2024honeycomb), synthesis route planning (liu2024multimodal), green chemistry (ruff2024using), and other areas. These studies demonstrate the advantages of LLMs in molecular representation and optimization.
In biology, LLMs are increasingly being used for tasks such as protein engineering (shen2024toursynbio), enzyme design (teukam2024integrating), and biological sequence analysis (feng2024large). By combining LLMs with vast amounts of biological data, they can more accurately predict interactions between biological molecules and significantly improve the efficiency of bioinformatics workflows. This has important implications for drug discovery and therapeutic protein design. The unique sequence generation and optimization capabilities of LLMs offer new possibilities for tackling combinatorial optimization challenges in biomacromolecular design.
Although applications in physics are relatively fewer, some emerging work has demonstrated the broad prospects of LLMs. pan2024quantum use multi-step prompt templates to prove that LLMs can perform complex analytical calculations in theoretical physics. Quantum computing algorithm design, physics simulation optimization, and computational methods in condensed matter.
5.3.3. Mechanics
MechAgents (ni2024mechagents) introduces a class of physics-inspired generative machine learning platforms that utilize multiple LLMs to solve mechanics problems, such as elasticity, by autonomously collaborating to write, execute, and self-correct code using finite element methods. Moreover, du2024large adopt LLMs in automatically discovering governing equations from data, utilizing the models’ generation and reasoning capabilities to iteratively refine candidate equations. This approach, tested on various nonlinear systems including the Burgers, Chafee–Infante, and Navier–Stokes equations, demonstrates a superior ability to uncover correct physical laws and shows better generalization on unseen data compared to other models. AutoTurb (zhang2024autoturb), on the other hand, adopts LLMs in an evolutionary framework to automatically design and search for turbulence model in computational fluid dynamics. Furthermore, buehler2023melm study LLM-based methods for forward and inverse mechanics problems including bio-inspired hierarchical honeycomb design, carbon nanotube mechanics, and protein unfolding.
Table 4. An Overview of Science Discovery Applications Utilizing Language Models Across Various Domains and Tasks.
Application Method Role of LLM Specific Problems or Tasks
General Scientific Equation Discovery Bilevel (ma2024llm) LLMaD Physical Scientific Discovery
LLM-SR (shojaee2024llm) LLMaD Scientific Equation Discovery
LLM4ED (du2024llm4ed) LLMaD Equation Discovery
Chemical ChatChemTS (ishida2024large) LLMaD Molecule Design
Debjyoti Bhattacharya, et al. (bhattacharya2024large) LLMaO Molecule Design
Agustinus Kristiadi, et al. (kristiadi2024sober) LLMaE Molecule Design
BoChemian (rankovic2023bochemian) LLMaE Chemical Reaction
Multi-modal MoLFormer (soares2024capturing) LLMaP Battery Electrolytes Formulation Design
CataLM (wang2024catalm) LLMaP Catalyst Design
Gavin Ye (ye2024novo) LLMaD Drug Design
DrugAssist (ye2023drugassist) LLMaO Drug Design
Biology MLDE (tran2024protein) LLMaP Protein Design
Prollama (lv2024prollama) Mixed Protein Design
X-LoRA (buehler2024x) LLMaP Protein Design
CodonBERT (li2023codonbert) LLMaP mRNA Design and Optimization
Revisiting-PLMs (hu2022exploring) LLMaP Protein Function Prediction
Mechanics MechAgents (ni2024mechagents) LLMaD Mechanics Design
MeLM (buehler2023melm) LLMaP, LLMaD Carbon Nanotubes and Proteins Design
Mengge Du, et al. (du2024large) LLMaD Nonlinear Dynamics Equation Discovery
AutoTurb (zhang2024autoturb) LLMaD Computational Fliud Dynamics
6. Challenges and Future Directions
6.1. Scalability
A primary limitation of LLMs in algorithm design is their scalability. LLMs operate within a fixed context window, which restricts the amount of information they can process simultaneously. This constraint is particularly problematic for complex algorithmic tasks that involve detailed specifications or large inputs. Furthermore, even with an adequate context window, LLMs often struggle to accurately comprehend and reason over long sequences of information and generate a long solution (chen2024design).
While the growing capabilities of LLMs and advancements in reasoning models (guo2025deepseek) have improved performance on tasks that can be mapped to their prior knowledge, this strength offers little advantage for novel algorithm design problems (liu2025interpretable). For instance, when used as optimizers, LLMs tend to perform well only on low-dimensional problems (yang2024large). Similarly, in the LLMaD paradigm, LLMs are often limited to generating individual components or heuristic functions rather than a complete algorithm framework (liu2024evolution).
To mitigate these scalability challenges, some researchers have proposed hybrid methods. For example, novikov2025alphaevolve introduce an approach that combines targeted code modifications using diff blocks with complete code regeneration, allowing for the evolution of more complex programs. Although related research has explored long-code generation (Ishibashi2024selforganized), these efforts typically prioritize code correctness and general software engineering over the specific challenge of algorithmic scalability.
6.2. Generalization
Another major challenge lies in generalization across different distributions and problem types. Existing methods often exhibit strong performance within their training distribution but fail to generalize to unseen instances (sim2025beyond). Designing a single heuristic that performs optimally across diverse instance distributions is inherently difficult (sim2025beyond; shi2025generalizable). For example, sim2025beyond reveal that algorithms designed on synthetic bin packing instances perform poorly on bin packing instances with real-world distributions. Similarly, shi2025generalizable show that an algorithm designed for TSP instances with a fixed size can hardly generalize to TSP instances with other sizes.
Several approaches aim to enhance cross-distribution generalization. shi2025generalizable adopt a meta-learning framework that promotes adaptation across distributions. zhang2025llm partition the overall problem class into subclasses based on instance features, enabling differentiated and automated heuristic design for each subclass. In contrast to designing a single optimal algorithm, EoH-S (liu2025eoh) proposes the concept of heuristic set design, which creates a portfolio of complementary algorithms to improve overall robustness across heterogeneous distributions.
Beyond distribution generalization, cross-problem generalization presents an even greater challenge, particularly for LLM-driven algorithm design. The majority of existing works focus on developing algorithms for a specific target task, and the resulting algorithm can rarely be applied to solve other problems (liu2024evolution). On the one hand, the algorithmic code implementation can not be used for solving other problems. On the other hand, the designed algorithm idea is either too high-level or problem-specific and thus lack cross-problem generalization. This contrasts with commonly used algorithmic frameworks, such as metaheuristic pipelines, which are often applicable to various problems despite differences in specific code implementations, as the core algorithmic idea remains the same (gendreau2010handbook). While there has been some discussion on designing general metaheuristics (pluhacek2023leveraging), this work has remained at a high level without comprehensive evaluation. Consequently, cross-problem algorithm design is a promising and challenging direction for future research.
6.3. Interpretability
Due to the black-box nature of LLMs, it is often difficult to trace how specific inputs influence the generated outputs. Moreover, we lack a deep understanding of the iterative algorithm design process itself (allen2023physics; liu2025large).
Recent studies have begun to address this challenge by analyzing the algorithmic landscape of LLM-driven design. For instance, liu2025large employ a simpler, interpretable model to approximate the behavior of LLMs when used as optimizers for multi-objective problems. To visualize the design process, van2025code propose Code Evolution Graphs (CEGs), which map the evolution of code using abstract syntax tree (AST) features. liu2025fitness introduce a graph-based representation where nodes signify algorithms and edges denote their evolutionary relationships, enabling a structural analysis of the design space. A related and crucial research direction involves developing appropriate similarity metrics to quantify the relationships between generated algorithms, which remains an active area of investigation (yao2024multi; liu2025fitness).
Other work has focused on establishing a theoretical basis for these methods. lee2025convergence propose to interpret the LLMaO procedure as a finite-state Markov chain, providing a potential avenue for formal analysis.
6.4. Efficiency
Efficiency is another major challenge in applying LLMs to algorithm design, encompassing both efficient iterative algorithm search and evaluation reduction.
While early methods relied on simple evolutionary search (liu2024evolution), more sophisticated strategies have been introduced to improve search efficiency. These include reflective reasoning (ye2024reevo), Monte Carlo Tree Search (zhengmonte), and ensemble methods (novikov2025alphaevolve). Recent attempts have also explored fine-tuning LLMs specifically for algorithm design, using both offline (liu2025fine) and online (surina2025algorithm; huang2025calm) approaches. For instance, liu2025fine constructed diverse algorithm pairs and employed Direct Preference Optimization (DPO) to offline fine-tune a smaller model. This specialized model demonstrated higher inference speeds and achieved results competitive with much larger models. In an online setting, huang2025calm utilized reinforcement learning to dynamically update the model during the algorithm search, leading to faster convergence and superior final algorithms within a fixed computational budget. However, these fine-tuned models are typically specialized for code generation on a single task. A key open question is how to accelerate the entire algorithm design process in a way that generalizes across diverse problem domains.
Furthermore, real-world algorithm design often requires computationally expensive evaluations. This expense arises from two factors: the large number of test instances required for robust assessment and the time-consuming nature of each individual evaluation, which may involve complex simulations or searches. Recent work has attempted to reduce this evaluation cost by enabling the LLM to predict an algorithm’s quality and terminate unpromising evaluations early (jawahar2024llm). Despite these efforts, there remains a need for principled approaches that can effectively balance the trade-off between exploration efficiency and performance reliability.
6.5. Benchmarking
Benchmarking is essential for ensuring fairness, reproducibility, and standardization across studies, facilitating both qualitative and quantitative comparisons. Despite rapid progress in LLM4AD, the field currently lacks standardized benchmarks for either algorithm design tasks or evaluation pipelines assessing LLM capabilities in this context.
While related benchmarks have been developed for mathematical reasoning (lin2024atg), planning (valmeekam2024planbench), and classical algorithmic tasks (velivckovic2022clrs), they do not fully capture the nuances of LLM-driven design. Recently, efforts have begun to develop benchmarks specifically for LLM-driven algorithm design. For example, sun2025co and chen2025heurigym introduce benchmark suites focused on combinatorial optimization. van2025blade developed a benchmark for designing heuristics for black-box optimization. Broadening the scope, liu2024llm4ad proposed a collection of tasks spanning diverse domains, from optimization and machine learning to scientific discovery.
Despite this progress, establishing rigorous and widely accepted benchmarks with standard settings remains a significant challenge. The development of datasets, unified evaluation protocols, and transparent reporting practices is crucial. Such standards will not only enhance reproducibility but also catalyze innovation in this emerging research area.
7. Conclusion
In this survey, we have provided a systematic review of the emerging field of algorithm design with large language models (LLM4AD). We categorized existing works into four paradigms based on the LLM’s primary role: LLM as Optimizer (LLMaO), LLM as Predictor (LLMaP), LLM as Extractor (LLMaE), and LLM as Designer (LLMaD). The LLMaO paradigm employs the LLM as a black-box optimizer to generate solutions, while LLMaP uses it as a surrogate model for prediction. LLMaE leverages the LLM to derive semantic features from unstructured data to inform an algorithm, and LLMaD directly generates algorithmic components or designs the entire algorithm. Furthermore, we have mapped works to the three core stages of the algorithm design, i.e., ideation, implementation, and evaluation, highlighting both progress and limitations. Finally, we summarized key application domains and identified critical open challenges, including scalability, generalization, interpretability, efficiency, and benchmarking, to help guide future research in this area.
8. Acknowledgments
The work described in this paper was supported by the Research Grants Council of the Hong Kong Special Administrative Region, China [GRF Project No. CityU11217325].